OpenStudy (anonymous):

What is the solution to the equation 2(m-8)=4(m+6)? m =

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

ur right:)

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

2(m - 8) = 4(m + 6) 2m - 16 = 4m + 24 2m - 4m = 24 + 16 -2m = 40 m = 40/-2 m = -20 yep :)